OpenStudy (anonymous):

Part B In a similar fall, Jim's rope stretches by only 30 cm. How many times his weight will the rope pull on him? Express your answer using two significant figures. F =

OpenStudy (anonymous):

@agent0smith @Xishem

OpenStudy (anonymous):

would i use the same numbers

OpenStudy (agent0smith):

Basically.. find his velocity (it seems he also fell 2 m) v^2 = 0^2 + 2(9.8)(2) so v again is 6.261m/s then find a: 0 = 6.261^2 + 2a(0.3)

OpenStudy (anonymous):

i got 65.31 m/s2 = a

OpenStudy (agent0smith):

a=-65.33 m/s^2 Then find it the same way... net F = 65.33m, so div by 9.8, 6.666 times his weight  So T = 7.7 W to 2 sig. figs

OpenStudy (anonymous):

would this a negative

OpenStudy (anonymous):

wouldnt the acceleration be -

OpenStudy (agent0smith):

Well that depends if you make up negative and down positive... the net force is UP, his weight is down, so choose your negatives and positives... up is easier to take as positive (ignore the negative value on my acceleration)

OpenStudy (agent0smith):

net force is + since i decided up is positive, it doesn't change this either way